# From Tokens to Thoughts: How LLMs and Humans Trade Compression for Meaning

### v4

@@ Abstract
Humans organize knowledge into compact categories through semantic compression by mapping diverse instances to abstract representations while preserving meaning (e.g., robin and blue jay are both birds; most birds can fly). These concepts reflect a trade-off between expressive fidelity and representational simplicity. Large Language Models (LLMs) demonstrate remarkable linguistic abilities, yet whether their internal representations strike a human-like trade-off between compression and semantic fidelity is unclear. We introduce a novel information-theoretic framework, drawing from Rate-Distortion Theory and the Information Bottleneck principle, to quantitatively compare these strategies. Analyzing token embeddings from a diverse suite of LLMs against seminal human categorization benchmarks, we uncover key divergences. While LLMs form broad conceptual categories that align with human judgment, they struggle to capture the fine-grained semantic distinctions crucial for human understanding. More fundamentally, LLMs demonstrate a strong bias towards aggressive statistical compression, whereas human conceptual systems appear to prioritize adaptive nuance and contextual richness, even if this results in lower compressional efficiency by our measures. These findings illuminate critical differences between current AI and human cognitive architectures, guiding pathways toward LLMs with more human-aligned conceptual representations.
1 Introduction: The Enigma of Meaning in Large Language Models
@@
The human capacity for concept formation is a cornerstone of intelligence, enabling us to manage information overload by deriving meaning from complex signals. We achieve this by identifying essential features and compressing experiences into cognitively tractable summaries (Murphy, 2004). This conceptual architecture, often hierarchical (e.g., a robin is a bird, an animal (Rosch et al., 1976)), is a powerful semantic compression: diverse instances are mapped to compact representations. Crucially, this process balances representational efficiency (compression) with the preservation of essential semantic fidelity (meaning), a trade-off fundamental to learning and understanding.
Large Language Models (LLMs) exhibit striking capabilities in processing and generating human language, performing tasks that often appear to require deep semantic understanding (Singh et al., 2024; Li et al., 2024). Despite this, a fundamental enigma persists: Do LLMs truly grasp concepts and meaning analogously to humans, or is their success primarily rooted in sophisticated statistical pattern matching over vast datasets? This question is particularly salient given the human ability to effortlessly distill extensive input into compact, meaningful concepts, a process governed by the inherent trade-off between informational compression and semantic fidelity (Tversky, 1977; Rosch, 1973b).
As the mental scaffolding of human cognition, concepts enable efficient interpretation, generalization from sparse data, and rich communication. For LLMs to transcend surface-level mimicry and achieve more human-like understanding, it is critical to investigate how their internal representations navigate the crucial trade-off between information compression and the preservation of semantic meaning. Do LLMs develop conceptual structures mirroring the efficiency and richness of human thought, or do they employ fundamentally different representational strategies?
To address this, we introduce a novel quantitative methodology rooted in information theory. We develop and apply a framework drawing from Rate-Distortion Theory (Shannon, 1948) and the Information Bottleneck principle (Tishby et al., 2000) to systematically compare how LLMs and human conceptual structures balance representational complexity (compression) with semantic fidelity. As a crucial human baseline, we leverage seminal datasets from cognitive psychology detailing human categorization (Rosch, 1973a, 1975; McCloskey and Glucksberg, 1978). A contribution of this work is the digitization and public release of these classic datasets, which offer benchmarks of high empirical rigor often exceeding modern crowdsourced alternatives. Our framework is tailored to dissect how these different systems navigate the compression-meaning trade-off.
Our comparative analysis across a diverse suite of LLMs reveals divergent representational strategies. While LLMs generally form broad conceptual categories aligned with human judgment, they often fail to capture the fine-grained semantic distinctions pivotal to human understanding. More critically, we uncover a stark contrast in priorities: LLMs exhibit a strong drive towards aggressive statistical compression, whereas human conceptual systems appear to favor adaptive nuance and contextual richness, even at a potential cost to sheer compressional efficiency by our measures. This divergence underscores fundamental differences and informs pathways for developing AI with more human-aligned conceptual understanding.
@@ 2 Research Questions and Scope
Advancing AI beyond pattern matching towards deeper semantic understanding hinges on whether LLMs develop conceptual structures analogous to human cognition. Human concepts efficiently balance semantic richness with cognitive manageability, a trade-off between meaning and informational compression. This paper investigates if and how LLMs replicate this fundamental balance.
Prior work has explored the conceptual landscape of LLMs, including their grasp of relational knowledge (Shani et al., 2023; Misra et al., 2021), methods for extracting interpretable concepts (Hoang-Xuan et al., 2024; Maeda et al., 2024), emergent representations via sparse activations (Li et al., 2024), embedding geometry concerning hierarchies (Park et al., 2024), and autoregressive concept prediction (Barrault et al., 2024). While insightful, these studies often lack a deep, quantitative comparison of the compression-meaning trade-off using an information-theoretic lens benchmarked against rich human cognitive data, or they may not ground concept definitions in established cognitive theory. Consequently, a rigorous comparative evaluation of how LLMs and humans balance representational efficiency with semantic fidelity remains a key open area. Separately, cognitive science has applied information theory to human concept learning (Imel and Zaslavsky, 2024; Tucker et al., 2025; Wolff, 2019; Sorscher et al., 2022), yet typically without connecting to modern AI models.
This work aims to bridge this gap by integrating cognitive psychology, information theory, and modern NLP. We pose three central research questions to guide our investigation:
  • [RQ1]: To what extent do concepts emergent in LLMs
    align with human-defined conceptual categories?
  • [RQ2]: Do LLMs and humans exhibit similar internal geometric structures
    within these concepts, especially concerning item typicality?
  • [RQ3]: How do humans and LLMs differ in their strategies for balancing representational compression with the preservation of semantic fidelity when forming concepts?
These three questions steer our investigation, which approaches each through the unifying lens of the information-theoretic framework detailed in Section 4. RQ1 begins by examining the alignment of broad conceptual categories, a key aspect of how information is compressed. RQ2 then delves into the finer-grained internal structures of these categories, probing the preservation of semantic nuances such as item typicality. Building on these analyses, RQ3 employs the full framework to comprehensively compare how LLMs and humans may divergently optimize the overall trade-off between compression and meaning. To ground these comparisons, we consistently utilize seminal human categorization datasets (Rosch, 1973a, 1975; McCloskey and Glucksberg, 1978) as empirical benchmarks. Our overarching aim is to use this comparative, information-theoretic approach not only to evaluate current LLMs but also to advance our understanding of efficient and meaningful representation in both artificial and natural intelligence.
@@ 3 Benchmarking Against Human Cognition
Empirically investigating the relationship between LLM representations and human conceptual structures requires two critical components: robust benchmarks of human categorization and a diverse selection of LLMs. This section details these components.
3.1 Human Conceptual Baselines: Empirical Data from Seminal Cognitive Science
Our comparison is anchored by data from seminal studies in cognitive psychology that mapped human categorization processes. These studies offer rich empirical evidence of how humans form concepts, judge category membership, and perceive typicality. Critically, unlike many modern crowdsourced datasets which can be noisy, these classic benchmarks were meticulously curated by cognitive science experts, reflecting deep cognitive patterns rather than superficial associations, and were grounded in then-advancing theories of conceptual structure. We focus on three influential works:
Rosch (1973): This foundational work by Rosch (1973a) explored semantic categories as part of the research program leading to prototype theory (Rosch, 1973c). This theory posits that categories are organized around “prototypical” members rather than strict, equally shared features. The dataset includes 48 items in eight common semantic categories (e.g., furniture, bird), with prototypicality rankings (e.g., ‘robin‘ as a typical bird, ‘bat‘ as atypical).
Rosch (1975): Building on prototype theory, Rosch (1975) further detailed how semantic categories are cognitively represented. This work provides extensive typicality ratings for a larger set of 552 items across ten categories (e.g., ‘orange‘ as a prototypical fruit, ‘squash‘ as less so).
McCloskey & Glucksberg (1978): McCloskey and Glucksberg (1978) investigated the “fuzzy“ boundaries of natural categories, showing that membership is often graded rather than absolute. Their data covers 449 items in 18 categories, with typicality scores and membership certainty ratings (e.g., ‘dress‘ is typical clothing, ‘bandaid‘ less so).
While originating from different research groups with distinct theoretical emphases, these datasets share rigorous experimental designs and provide data on both category assignments and item typicality. We aggregated data from these studies, creating a unified benchmark of 1,049 items across 34 categories. This aggregated dataset, which we have digitized and make publicly available (see Appendix A.1), offers a crucial, high-fidelity empirical foundation for evaluating the human-likeness of computational models and we encourage its use for future research.
Appendix  A.6 complements this benchmark by reporting WordNet-based polysemy statistics for every item. These figures characterize the lexical ambiguity of the stimuli and frame the subsequent embedding analyses. The analysis shows that polysemy is very rare in our data and thus cannot explain the findings we report later in the paper.
@@ 3.2 Large Language Models Under Study
We include a diverse array of LLMs to assess how conceptual representation might vary with computational architecture and scale. This selection covers prevalent architectural paradigms (encoder-only, decoder-only) and a wide spectrum of model sizes, from 300 million to 72 billion parameters.
Our analysis features encoder-only models from the BERT family (e.g., BERT-Large (Devlin et al., 2019; He et al., 2020; Zhuang et al., 2021)). The majority are decoder-only autoregressive models, including: six Llama family models (1B to 70B, e.g., Llama 3.1 70B (Touvron et al., 2023a, b; Grattafiori et al., 2024)); five Gemma family models (2B to 27B (Team et al., 2024, 2025)); thirteen Qwen family models (0.5B to 72B (Bai et al., 2023; Yang et al., 2024)); four Phi family models (e.g., Phi-4 (Javaheripi et al., 2023; Abdin et al., 2024; Abouelenin et al., 2025)); a Mistral 7B model (Karamcheti et al., 2021); two GPT-2 models of different scales (Radford et al., 2019); and checkpoints from the OLMo-7B model (Groeneveld et al., 2024). Other, less common architectures include a single encoder-decoder T5 model (Raffel et al., 2020), two text encoders from CLIP ViT (ViT-B/32 and ViT-B/16) (Radford et al., 2021), and the classic encoder models GloVe (Pennington et al., 2014) and Word2Vec (Mikolov et al., 2013a, b). Note: GloVe and Word2Vec provide static embeddings only, and thus their analysis is constrained. Appendix A.2 provides a comprehensive list of all model variants, identifiers, and architectural details.
For each LLM, we extract static, token-level embeddings from its input embedding layer (the ‘E‘ matrix). This choice aligns our analysis with the context-free nature of stimuli typical in human categorization experiments, ensuring a comparable representational basis. These embeddings form the foundation for deriving LLM-generated conceptual clusters in our subsequent analyses.
To complement the analysis, we extend our methodology beyond static embeddings to contextual representations. For each LLM, we extract token-level embeddings using a set of carefully controlled prompts (see Appendix A.3), enabling a principled basis for directly comparing static token representations to their contextualized counterparts. This allows for a similar assessment while mitigating confounding issues such as prompt length and token position. We found minimal variation in results across prompt templates and pooling methods. Appendix A.3 provides a detailed exploration.
Finally, tokenizer-overlap analysis (Appendix A.7) shows substantial lexical sharing across model families; neither token count nor tokenization overlap correlates with AMI, confirming that tokenizer effects are negligible and thus cannot explain our findings.
@@ 4 A Framework for Comparing Compression and Meaning
To understand how LLMs and human cognition grapple with the fundamental challenge of representing meaning, we introduce an information-theoretic framework. This framework is designed to analyze the critical trade-off, or tension, between compressing information into efficient representations and preserving the rich semantic fidelity essential for true understanding. Drawing upon core principles from Rate-Distortion Theory (RDT) (Shannon, 1948) and the Information Bottleneck (IB) principle (Tishby et al., 2000), our approach provides a cohesive lens for addressing all three of our research questions. Our investigation progresses by first exploring distinct facets of this trade-off related to representational compactness and semantic preservation, before synthesizing these insights to evaluate the overall efficiency of conceptual representation. Our research questions, viewed through this progressive information-theoretic perspective, are approached as follows:
  • [RQ1] Probing Representational Compactness via Categorical Alignment: We begin by examining how information is condensed into categorical structures. Both human categorization and LLM-derived clustering simplify diverse items $X$ into structured groups $C$. For RQ1, we assess alignment between model-based clusters ($C_{LLM}$) and human categories ($C_{Human}$) by quantifying shared information (e.g., via Adjusted Mutual Information), offering an initial view on how similarly compactness is achieved. The principles of efficient input representation here relate to the “Complexity” aspect of our framework.
  • [RQ2] Probing Semantic Preservation via Internal Structure: Next, we assess how well meaning is preserved within these compressed representations. An effective system must retain crucial semantic nuances. For RQ2, we investigate this by correlating LLM-internal measures of item centrality with human typicality judgments, probing how faithfully fine-grained semantic information is represented, that is, can LLMs capture the internal structure of $C_{Human}$? This relates to the “Distortion” (or fidelity) aspect of our framework.
  • [RQ3] Evaluating the Integrated Trade-off for Total Representational Efficiency: Finally, having explored compactness and preservation, we leverage our full framework. RQ3 employs a unified objective function, $\mathcal{L}$ (detailed below), to quantitatively assess the total efficiency with which LLMs and human systems navigate this fundamental trade-off.
The following subsections detail the theoretical underpinnings of this framework.
4.1 Theoretical Underpinnings: Rate-Distortion Theory and the Information Bottleneck
To rigorously formalize the balance between representational compactness and preserved meaning, we draw upon information theory. Rate-Distortion Theory (RDT) (Shannon, 1948) provides the foundational language. RDT quantifies the minimal “rate” $R$ (representational complexity) needed to represent a source $X$ as $C$, subject to a maximum “distortion” $D$ (fidelity loss). The goal is often to optimize $R+\lambda D$, offering a principled evaluation of representational efficiency.
The Information Bottleneck (IB) principle (Tishby et al., 2000) is a related approach. IB seeks a compressed representation $C$ of an input $X$ that maximizes information about a relevant variable $Y$ while minimizing $I(X;C)$, the mutual information $C$ retains about $X$ (the bottleneck’s “cost”). This is typically framed as minimizing $I(X;C)-\beta I(C;Y)$.
Our analytical framework directly applies RDT’s core idea of balancing rate and distortion. We formulate an objective function, $\mathcal{L}$, designed to explicitly balance a complexity term (analogous to RDT’s rate), which quantifies the informational cost of representing items $X$ through their conceptual clusters $C$, and a distortion term (analogous to RDT’s $D$), which measures semantic information lost or obscured within these clusters. Our complexity term, incorporating $I(X;C)$, resonates with the IB principle. However, our distortion term directly measures intra-cluster semantic fidelity loss (specifically, the variance of item embeddings relative to their cluster centroids), differing from canonical IB formulations where distortion is often implicitly tied to an external relevance variable $Y$. This direct approach allows us to evaluate how any given clustering $C$, whether derived from human cognitive data or LLM embeddings, intrinsically balances its own structural compactness and the meaningfulness of its components with respect to the original data $X$.
4.2 The $\mathcal{L}$ Objective: Balancing Representational Complexity and Semantic Distortion
Building on these information-theoretic foundations, this section formally defines the two key components of our framework–Complexity and Distortion. These components allow us to quantitatively address the aspects of representational compactness (core to [RQ1]) and semantic preservation (central to [RQ2]) that were introduced earlier. We then combine these into a unified objective function, $\mathcal{L}$, designed to evaluate the overall efficiency of the compression-meaning trade-off, which is the primary focus of [RQ3]. The $\mathcal{L}$ function evaluates the efficiency of the conceptual clusters $C$ derived from items $X$ (e.g., token embeddings):

### v5

@@ Abstract
Humans organize knowledge into compact categories that balance compression with semantic meaning preservation. Large Language Models (LLMs) demonstrate striking linguistic abilities, yet whether they achieve this same balance remains unclear. We apply the Information Bottleneck principle to quantitatively compare how LLMs and humans navigate this compression-meaning trade-off. Analyzing embeddings from 40+ LLMs against classic human categorization benchmarks (Rosch, 1973a; 1975; McCloskey & Glucksberg, 1978), we uncover three key findings. First, LLMs broadly align with human categories but miss fine-grained semantic distinctions crucial for human understanding. Second, LLMs demonstrate aggressive statistical compression, achieving “optimal” information-theoretic efficiency, while humans prioritize contextual richness and adaptive flexibility. Third, encoder models surprisingly outperform decoder models in human alignment, suggesting that generation and understanding rely on distinct mechanisms in current architectures. In addition, training dynamics analysis reveals that conceptual structure develops in distinct phases: rapid initial formation followed by architectural reorganization, with semantic processing migrating from deeper to mid-network layers as models discover more efficient encoding. These divergent strategies, where LLMs optimize for compression and humans for adaptive utility, reveal fundamental differences between artificial and biological intelligence, guiding development toward more human-aligned AI.
1 The Enigma of Meaning in Large Language Models
@@
Humans excel at organizing knowledge into concepts which are compact categories that achieve remarkable compression while preserving essential meaning (Murphy, 2004). A single word like “bird” compresses information about thousands of species, yet maintains critical semantic properties (can fly, has feathers, lays eggs). This hierarchical organization (robin $\rightarrow$ bird $\rightarrow$ animal; Rosch et al. 1976) represents a fundamental cognitive achievement: balancing efficiency with semantic fidelity.
Large Language Models (LLMs) demonstrate striking linguistic capabilities that suggest semantic understanding (Singh et al., 2024; Li et al., 2024). Yet, a critical question remains unanswered: Do LLMs navigate the compression-meaning trade-off similarly to humans, or do they employ fundamentally different representational strategies? This question matters because true understanding, which goes beyond surface-level mimicry, requires representations that balance statistical efficiency with semantic richness (Tversky, 1977; Rosch, 1973b).
To address this question, we apply Rate-Distortion Theory (Shannon, 1948) and Information Bottleneck principles (Tishby et al., 2000) to systematically compare LLM and human conceptual structures. We digitize and release seminal cognitive psychology datasets (Rosch, 1973b; 1975; McCloskey & Glucksberg, 1978), which are foundational studies that shaped our understanding of human categorization but were previously unavailable in a machine-readable form. These benchmarks, comprising 1,049 items across 34 categories with both membership and typicality ratings, offer unprecedented empirical grounding for evaluating whether LLMs truly understand concepts as humans do. It also offers much better quality data than the current crowdsourcing paradigm.
Analyzing embeddings from 40+ diverse LLMs against these benchmarks, we uncover a fundamental divergence: LLMs and humans employ different strategies when balancing compression with meaning. While LLMs achieve broad categorical alignment with human judgment, they optimize for aggressive statistical compression at the expense of semantic nuance. Humans maintain “inefficient” representations that preserve rich, multidimensional structure essential for flexible reasoning.
This divergence manifests across three dimensions. First, LLMs capture categorical boundaries but miss fine-grained semantic distinctions like item typicality central to human understanding. Second, our information-theoretic analysis reveals LLMs achieve mathematically “optimal” compression-distortion trade-offs, while human categories appear suboptimal. Third, encoder models surprisingly outperform decoder models in human alignment despite smaller scales, indicating that understanding and generation may require fundamentally different representational strategies.
Through analysis of OLMo-7B across 57 training checkpoints, we further uncover how these strategies emerge during learning: conceptual structure develops via rapid initial formation followed by architectural reorganization, with semantic processing migrating from deep to mid-network layers as models discover increasingly efficient encodings.
These findings challenge the assumption that statistical optimality equals understanding. The apparent “inefficiency” of human concepts may reflect optimization for adaptive flexibility. Our framework and newly-digitized benchmarks provide essential tools for monitoring this critical balance, guiding development toward AI systems that achieve not just compression, but comprehension.
@@ 2 Research Questions and Scope
Prior work has explored LLM conceptual representations through multiple lenses: relational knowledge (Shani et al., 2023; Misra et al., 2021), interpretable concept extraction (Hoang-Xuan et al., 2024; Maeda et al., 2024), sparse activation patterns (Li et al., 2024), and embedding geometry including hierarchical structures (Park et al., 2024). While insightful, these studies often lack deep, quantitative comparison of the compression-meaning trade-off using information theory against rich human cognitive benchmarks.
Separately, cognitive science has applied information theory to human concept learning (Imel & Zaslavsky, 2024; Tucker et al., 2025; Zaslavsky et al., 2018; Sorscher et al., 2022). For example, Zaslavsky et al. (2018) developed an Information Bottleneck framework for color naming efficiency, later extended to animal taxonomies (Zaslavsky et al., 2019). Yet these cognitive studies typically proceed without connecting to modern LLMs, and tend to focus on a specific domain.
These two streams, LLM conceptual analysis and cognitive information theory, rarely intersect. We bridge this gap through rigorous comparison of how LLMs and humans navigate the compression-meaning trade-off, grounding our analysis in established cognitive benchmarks. This leads to three research questions:
  • [RQ1] To what extent do LLM-emergent concepts align with human-defined categories?
  • [RQ2] Do LLMs exhibit human-like internal structure, particularly item typicality?
  • [RQ3] How do humans and LLMs differ when balancing compression with semantic fidelity?
Our framework approaches each RQ through a unified lens. [RQ1] examines the categorical alignment, or how information is compressed into discrete groups. [RQ2] probes internal structure, which means how semantic meaning is preserved within categories. [RQ3] employs our full $\mathcal{L}$ objective to evaluate the integrated trade-off. This progression from compression to preservation to their balance mirrors the fundamental challenge both systems face: creating representations that are simultaneously efficient and meaningful.
@@ 3 Benchmarking Against Human Cognition
Investigating LLM-human conceptual alignment requires robust benchmarks and diverse models. This section details both components.
3.1 Human Baselines: Empirical Data from Seminal Cognitive Science
We draw on three foundational studies that shaped our understanding of human categorization. Unlike many noisy modern crowdsourced datasets, these classic benchmarks were carefully curated by experts, capturing deep cognitive patterns. We focus on three influential works:
Rosch (1973): This foundational work (Rosch, 1973a) explored semantic categories as part of the research program leading to prototype theory (Rosch, 1973b). The theory posits that categories organize around ”prototypical” members rather than strict, equally shared features. The dataset includes 48 items in eight common semantic categories (e.g., furniture, bird), with prototypicality rankings (e.g., ‘robin’ as typical bird, ‘bat’ as atypical).
Rosch (1975): Building on prototype theory, Rosch (1975) further detailed how semantic categories are cognitively represented. This work provides typicality ratings for a larger set of 552 items across ten categories (e.g., ‘orange‘ as a prototypical fruit, ‘squash‘ as less so).
McCloskey & Glucksberg (1978): Investigated the ”fuzzy” boundaries of natural categories, showing membership is graded rather than absolute (McCloskey & Glucksberg, 1978). Covers 449 items in 18 categories with both typicality scores and membership certainty ratings (e.g., ‘dress’ is typical clothing, ‘bandaid’ less so).
While originating from different researchers, these datasets share rigorous experimental designs and provide data on both category assignments and item typicality. We aggregated data from these studies, creating a unified benchmark of 1,049 items across 34 categories. This data, which we have digitized and made publicly available (Appendix B.1), offers a high-quality empirical foundation for evaluating the human-likeness of LLMs. [^1] Appendix B.6 shows that polysemy is rare in our data and cannot account for our findings.
@@ 3.2 Large Language Models Under Study
We analyze 40+ diverse LLMs spanning multiple architectures and scales (300M to 72B parameters) to understand how conceptual representation varies across model design choices.
Model Selection. Our study encompasses three architectural paradigms. Encoder models include the BERT family (Devlin et al., 2019; He et al., 2020; Zhuang et al., 2021) and CLIP ViT text encoders (Radford et al., 2021). Decoder models form the majority of our analysis: the Llama family (1B-70B; Touvron et al., 2023a; b; Grattafiori et al., 2024), Gemma variants (2B-27B; Team et al., 2024; 2025), Qwen models (0.5B-72B; Bai et al., 2023; Yang et al., 2024), Phi series (Javaheripi et al., 2023; Abdin et al., 2024; Abouelenin et al., 2025), Mistral-7B (Karamcheti et al., 2021), GPT-2 (Radford et al., 2019), and OLMo-7B (Groeneveld et al., 2024). We also include classic embeddings Word2Vec (Mikolov et al., 2013a; b) and GloVe (Pennington et al., 2014) as baselines.
This diverse selection enables us to disentangle effects of architecture (encoder vs. decoder), scale (300M to 72B), and training objectives (understanding vs. generation). Complete model specifications appear in Appendix B.2.
Embedding Extraction. We extract representations at two levels to capture different aspects of conceptual knowledge: (1) static embeddings from input layers (E matrix), capturing context-free lexical knowledge directly comparable to isolated words in human categorization experiments; and (2) contextual embeddings from hidden layers using controlled prompts, revealing how context shapes conceptual structure across network depth.
This dual approach allows us to trace how concepts emerge from basic lexical knowledge to contextualized understanding. Critically, our results prove robust to prompt templates and pooling strategies (Appendix B.3). Moreover, despite substantial vocabulary overlap between model families, neither token count nor tokenization patterns correlate with our results (Appendix B.7).
@@ 4 A Framework for Comparing Compression and Meaning
To quantitatively compare how LLMs and humans navigate the fundamental tension between compact representation and semantic richness, we develop a framework that captures both aspects of conceptual organization. Our approach draws on Rate-Distortion Theory (Shannon, 1948) and the Information Bottleneck principle (Tishby et al., 2000), adapting them to measure the quality of conceptual systems.
4.1 Theoretical Foundations
Human concepts achieve remarkable efficiency: the word ”bird” compresses knowledge about thousands of species into a single category, yet preserves critical semantic information (can fly, has feathers, lays eggs). This reflects a fundamental trade-off that any conceptual system must navigate:
  • Compression: Grouping diverse items into manageable categories (fewer bits needed)
  • Meaning Preservation: Maintaining semantic coherence within groups
Rate-Distortion Theory (RDT; Shannon, 1948) formalizes this trade-off for lossy compression. Given data $X$ and compressed representation $\hat{X}$, RDT seeks encodings that minimize:
[…]
where $R$ is the rate (bits required), $D$ is distortion (information lost), and $\lambda$ controls their trade-off.
Information Bottleneck (IB; Tishby et al., 2000) extends this by compressing $X$ into $Z$ while preserving information about relevant variable $Y$:
$\min I(X;Z)-\beta I(Z;Y)$ (2)
Our adaptation: For conceptual representation, we lack an external relevance variable $Y$. Instead,“relevance” becomes internal semantic coherence, which means how well categories preserve within-group similarity. We thus combine RDT’s geometric distortion with IB’s information-theoretic compression, yielding our framework where clustering $C$ represents items $X$ by minimizing both the information needed to specify items (compression) and the semantic spread within clusters (distortion).
4.2 The $\mathcal{L}$ Objective: Quantifying the Trade-off
We formalize how clustering $C$ represents items $X$ through an objective that combines information-theoretic compression with geometric coherence: